OpenStudy (angel_kitty12):

P(B)=2/3 P(A)=2/5 P(A or B)=11/20 Find P(A|B)

OpenStudy (freckles):

do you know another way to write P(A|B)?

OpenStudy (freckles):

Also do you recall: \[P(A \text{ or } B)=P(A)+P(B)-P(A \text{ and } B) \\ \text{ and so } P(A \text{ and } B)=P(A)+P(B)-P(A \text{ or } B)\]

OpenStudy (angel_kitty12):

P(A|B)= P(A)•P(B)-P(A and B)

OpenStudy (angel_kitty12):

I believe so

OpenStudy (freckles):

shouldn't this be it: \[P(A|B)=\frac{P(A \text{ and } B)}{P(B)}\]

OpenStudy (angel_kitty12):

Oh that as well yes

OpenStudy (angel_kitty12):

I'm a bit confused with the formulas

OpenStudy (freckles):

never seen the first one you mentioned are you sure that is right?

OpenStudy (freckles):

I definitely seen the one I mentioned

OpenStudy (angel_kitty12):

I don't think I'm right but that is a formula for dependent overlap but I think I mixed my formulas up

OpenStudy (angel_kitty12):

No wait IM WRONG I WROTE IT WRONG DISREGARD 😅😅 sorry

OpenStudy (freckles):

k well try using the one I mentioned try using both the ones I mentioned

OpenStudy (freckles):

like you find P(A and B) first using the first one I mentioned

OpenStudy (freckles):

then you can plug in that result a long with the number given for P(B) to find P(A|B)

OpenStudy (freckles):

so first: do you think you can find P(A and B) using: \[P(A \text{ and } B)=P(A)+P(B)-P(A \text{ or } B)?\]

OpenStudy (angel_kitty12):

Okay so the fraction forms would go in as 2/3+2/5-11/20 correct?

OpenStudy (freckles):

correct which equals

OpenStudy (angel_kitty12):

I'm a bit confused here though. Do I add the fractions even thought they do not have the same denominator?

OpenStudy (freckles):

to add/subtract fractions you need to find a common denominator first

OpenStudy (angel_kitty12):

Okay that's what confused me. Let me try..

OpenStudy (angel_kitty12):

31/60

OpenStudy (freckles):

k know use the second formula we mentioned you just found P(A and B) and you were already given P(B) plug into the formula for P(A|B)

OpenStudy (angel_kitty12):

I cannot see the second formula you have given

OpenStudy (freckles):

\[P(A|B)=\frac{P(A \text{ and } B)}{P(B)}\] you cannot see this one?

OpenStudy (angel_kitty12):

Wait yes I can now never mind

OpenStudy (angel_kitty12):

So it would kinda look like X=31/60/2/3 correct?

OpenStudy (freckles):

\[P(A|B)=\frac{\frac{31}{60}}{\frac{2}{3}} \text{ or can be written as } \frac{31}{60} \cdot \frac{3}{2}\]

OpenStudy (angel_kitty12):

So the answer would be 93/120 or 0.775 correct?

OpenStudy (freckles):

yes 93/120 can be reduced \[\frac{31}{60 } \cdot \frac{3}{2} =\frac{3}{60 } \cdot \frac{31}{2} =\frac{1}{20} \cdot \frac{31}{2}=\frac{31}{40}\] so 31/40 or .775

OpenStudy (angel_kitty12):

Or 31/40 if simplified

OpenStudy (angel_kitty12):

Okay I understand now. Thank you so much!